OpenStudy (anonymous):

A new branch in its first year opened new accounts at a rate of 40 every 100 customers calls. The total number of new accounts opened in the year was 10000. How many customer calls were made?

OpenStudy (aum):

Direct proportion: 40 / 100 = 10,000 / x OR 100 / 40 = x / 10000 Solve for x.